OpenStudy (anonymous):

Find the slope of g(x) with g(10)=52; g(6)=42

OpenStudy (jdoe0001):

\(\bf g(10)=52\to \begin{cases} x=10\\ \bf y=52 \end{cases}\to (10,52) \\ \quad \\ g(6)=42\to \begin{cases} x=6\\ \bf y=42 \end{cases}\to (6,42) \\ \quad \\ \textit{so the line passes through }(10,52)\ and\ (6,42) \\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 10}}\quad ,&{\color{blue}{ 52}})\quad &({\color{red}{ 6}}\quad ,&{\color{blue}{ 42}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)